OpenStudy (anonymous):

Jasmine decides to solve this system of equations using the substitution method. 3x + y = 18 4x + 2y = 10 These are the first two steps of the method that Jasmine will use. Step 1: Solve for y in the first equation. Step 2: Substitute the result from Step 1 into the second equation. Which equation should be solved in Step 3 using Jasmine's method?

OpenStudy (anonymous):

a. 4x + 18 + 3x = 10 b. 4x + 18 - 3x = 10 c. 4x + 36 - 3x = 10 d. 4x + 36 + 6x = 10 e. 4x + 36 - 6x = 10

OpenStudy (***[isuru]***):

 hope this will help ya!!